If it were your goal to isolate (-)-phenylsuccinic acid instead of (+)-phenylsuccinic acid, suggest two modifications that yoExperiment:

  • Lab goals
    • Isolate an enantiopure sample of phenylsuccinic acid from a racemic mixture
    • Assess the enantiopurity of the recovered material using polarimetry experiment
  • Procedure
    • 1. Selectively preiciptate the (+) enantiomer of phenylsuccinic acid using L(–) proline resolving agent
    • 2. Remove the (–) proline (resolving agent) from the precipitate by washing with acid solution (HCl)
    • 3. Determine the enantiomeric excess of (+)-phenylsuccinic acid in the precipitate using polarimetry
    • 4. Collect your enantiomerically enriched product in the (+)-phenylsuccinic acid container
Answer #1

The resolution of racemic phenylsuccinic acid can be achieved by salt formation with naturally occuring (S)-proline. Opticall

Experimental details:

Dissolve 1.94 g (0.01 mol) of racemic phenylsuccinic acid in 50 mL of 2-propanol in a clean, dry 250-mL round-bottom flask. Add 1.15 g (0.01 mol) of finely ground (–)-proline. Mix the contents of the flask for 2–3 min with a stirring rod and then reflux for 30 min. During the reflux, as all of the proline dissolves, the (+)-bis proline salt usually begins to precipitate out of solution. Air cool the flask for 15–20 min to approximately 30 °C. If the (+)-salt has not crystallized at this point, vigorous stirring will initiate crystallization. Filter and wash the solid twice with 15 mL of acetone. Let the aspirator run for several minutes after the second acetone wash and then dry the salt between several pieces of filter paper. To isolate the (+)- phenylsuccinic acid in one lab period, the salt is hydrolyzed even though it is not completely dry. If desired, one may take the melting point and optical rotation of the salt (mp 160–161 °C, [α]D = +13.5° (c 1.00, methanol), by spreading the crystals on a clean filter paper and allowing it to thoroughly dry (ca. 60 min). Add all of the (+)-salt to 10 mL of 6M HCl that is being cooled in an ice water bath. Stir this mixture for 5 min. Filter and wash the solid twice with 15 mL of cold water. After that the solid is allowed to air dry for at least 24 hours. [α]D = +151° (c 4.0, acetone). Recrystallization from water will yield optically pure (+)-phenylsuccinic acid, with a melting point of 185–186 °C and [α]D = +171° (c 4.0, acetone) To isolate (–)-phenylsuccinic acid, evaporate the filtrate after filtering out the (+)-salt. To simply demonstrate that (–)-phenylsuccinic acid was in the filtrate, stir the remaining residue with 10 mL of 6M HCl for 10 min. Filtration of this mixture will yield optically active (–)-phenylsuccinic acid: 0.61 g, mp 169–171 °C, [α]D = 67° (c 4.0, acetone). Alternatively, if optically pure (–)-phenylsuccinic acid is desired, the residue is dissolved in 40 mL of absolute ethanol. Add 0.39 g of (–)-proline and stir at room temperature for 30 min. Evaporate off the ethanol to isolate the (–)-monoproline salt. Add 20 mL of acetone to the (–)-salt and stir for 2–3 min. Filter, and then dry the salt between several pieces of filter paper. (If the salt is allowed to dry thoroughly, typical results of the (–)-salt are 1.14 g, mp = 142–144 °C, [α]D = 96° (c 1.00, methanol)). Add all of the (–)-salt to 10 mL of 6M HCl and stir at 0 °C for 5 min. Filter and wash the solid twice with 15 mL of water. Let the (–)-phenylsuccinic acid air dry: 0.57 g, mp 176–179 °C, [α]D = 160° (c 4.0, acetone). Recrystallization from water will yield optically pure (–)-phenylsuccinic acid, 0.38 g (39%), mp 185– 186 °C, and [α]D = 170° (c 4.0, acetone).